# Conditional-$t^{3}$VAE: Equitable Latent Space Allocation for Fair Generation

### v1

@@ E.1 Datasets
    Test set 1744 5099 4149 2882 2523 2384 1977 2019 1660 1595
  • CIFAR100-LT : This balanced dataset comprised of colored images of 100 classes of different natural objects. The images are of size $32\times 32\times 3$ presents high variability with 100 fine-grained classes. In the most extreme imposed imbalance setting, tail classes may contain as few as five examples. Though the images remain low-resolution, this setting stresses the models’ robustness in the face of sparse data and various categories.
  • CelebA : This dataset enables evaluation in real-world class-imbalance scenarios using attribute labels (e.g., Mustache, Young). It also introduces the complexity of higher resolution images ($178\times 218\times 3$), testing the scalability of our models. We select four binary attributes (Mustache, Young, Male, Smiling) for evaluation, with test sets balanced by down-sampling the larger class. In Table 3, we provide the number of images for each chosen attribute in this dataset.
    Table 3: The Number of images in the CelebA dataset for the train and test sets for the Mustache, Young, Male and Smiling attributes.
@@
    Test set 722 19190 15114 4848 7715 12247 9987 9975
E.2 Data Preprocessing
  • SVHN-LT Dataset For both training and testing, we crop each class to the minimum number of samples available across all classes. The only data augmentation applied is a random horizontal flip with 50% probability.
  • CIFAR100-LT Dataset We use the dataset in its entirety without class filtering. As with SVHN-LT, we apply a random horizontal flip with 50% probability for data augmentation.
  • CelebA Dataset We use the full dataset and apply a center crop of $160\times 160$ to each image and then resize to $128\times 128$. A random horizontal flip with 50% probability is also applied.
E.3 Model Architecture
@@ E.4 Training Details
Appendix F Hyperparameter Tuning
@@ F.1.2 On the CIFAR100-LT dataset
F.1.3 On the CelebA dataset
For the CelebA dataset, we focus only on $\beta$ optimization for Student’s t models. As for the Gaussian-based models, we set $\beta=0.1$, because as shown in Table 1 the optimization of the $\beta$ hyperparameter has marginal impact on the FID for the CelebA dataset, which is not the case for the SVHN-LT and CIFAR100-LT dataset. Hence, we did not deem necessary to perform a hyper-parameter optimization over $\beta$ for the Gaussian based models. Nevertheless, from Figure 11 we see that $\beta$ has marginal effect on the Student’s t based models too, likely due to the low intra-class variability.
@@ F.1.3 On the CelebA dataset
(a) CelebA - Male
(b) CelebA - Smiling
@@
F.2 $\nu$ Optimization
Table 4 shows results from tuning the degrees of freedom parameter $\nu$ in C-$t^{3}$VAE on both SVHN-LT and CIFAR100-LT across all imbalance ratios, using the optimal $\beta$ from the previous study. On average, $\nu=10$ performs well, consistent with prior work [10]. However, performance can be further improved by selecting $\nu$ within the range $[2.5,20]$. Still, no major influence of hyper-parameter $\nu$ on the generative FID can be observed similarly to what was observed by [10] for the reconstruction FID.
@@ F.2 $\nu$ Optimization
Table 4: Variability of the FID as a function of the standard deviation $\nu$ for the C-$t^{3}$VAE model on the SVHN-LT and CIFAR100-LT datasets.
F.3 $\tau$ Optimization
In this section, we evaluate the effect of the $\tau$ parameter on the SVHN-LT, CIFAR100-LT and CelebA datasets for all imbalance ratios while setting $\beta$ and $\nu$ to their previously optimized values. As shown in Figure 13, the optimal $\tau$ for SVHN-LT aligns closely with our theoretical prediction. In contrast, CIFAR100-LT consistently benefits from a larger $\tau=0.4$, yielding improved FID across all imbalance settings and outperforming C-VAE. On CelebA, $\tau$ has minimal impact and the most likely value is $\tau\approx 0.3$. This value is between the value observed in for the SVHN-LT and CIFAR100-LT, as is mostly liked to the complexity and the entropy of the dataset where the more complex the data is the higher the value of $\tau$ is required.
@@ F.3 $\tau$ Optimization
(a) CelebA Young
(b) CelebA Male
(c) CelebA Smiling
@@
Appendix G Per-Class Evaluation

### v2

@@ D.1 Datasets
    Test set 1744 05099 04149 2882 2523 2384 1977 2019 1660 1595
    For both training and testing, we crop each class to the minimum number of samples available across all classes. The only data augmentation applied is a random horizontal flip with 50% probability.
  • CIFAR100-LT : The CIFAR100 dataset (Krizhevsky, 2009) consists of 60,00060{,}000 color images of size $32\times 32$ pixels, evenly distributed across $100$ object categories. Each category contains $600$ images, split into $500$ training samples and $100$ test samples. Figure 7 shows sample images after preprocessing.
    Figure 7: Sample images from the CIFAR100 dataset (Krizhevsky, 2009).
    We use the dataset in its entirety without class filtering. As with SVHN-LT, we apply a random horizontal flip with 50% probability for data augmentation.
  • CelebA : The CelebFaces Attributes Dataset (CelebA) (Liu et al., 2015) contains 202,599 color images of celebrity faces at a resolution of $178\times 218$. Each image is annotated with 40 binary facial attributes, such as Mustache, Smiling, and Young. This dataset exhibits significant variability in pose, expression, and illumination while providing high resolution images. We preprocess the images of this dataset by cropping the central region to $160\times 160$ and resizing to $128\times 128$ using bilinear interpolation. Figure 8 illustrates sample images after preprocessing. Also, we adhere to the original training, validation, and test splits provided by the dataset.
    Figure 8: Sample images from the CelebA dataset after preprocessing (Liu et al., 2015).
@@
    Test set 0722 019190 015114 04848 07715 12247 09987 09975
    Table 3: Number of images in the CelebA dataset for the train and test sets for the Mustache, Young, Male and Smiling attributes.
    Given CelebA’s multi-attribute structure, we select four binary attributes—Mustache, Young, Male, and Smiling—for training and evaluation, treating each attribute and its negation as distinct classes. These attributes were chosen due to their varying imbalance ratios, defined as the frequency of the majority class divided by the minority. The resulting ratios are 25 for Mustache, 3.5 for Young, 1.4 for Male, and 1 for Smiling. This enables the study of imbalance effects in generative modeling. We also balance the test sets by downsampling the larger class for each attribute. In Table 3 we report the number of images per selected attribute in the dataset.
D.2 Model Architecture
@@ D.3 Training Details
Appendix E Hyperparameter Tuning
@@ E.1.2 On the CIFAR100-LT dataset
E.1.3 On the CelebA dataset
For CelebA, we optimize $\beta$ exclusively for Student’s $t$ models, setting $\beta=0.1$ for Gaussian variants. Table 1 indicates that $\beta$ has minimal impact on CelebA’s FID, unlike the trends seen in SVHN-LT and CIFAR100-LT. Therefore, extensive tuning for Gaussian models is omitted. Figure 11 suggests that Student’s t models share this robustness to $\beta$ variations, likely driven by the dataset’s limited intra-class variability.
@@ E.1.3 On the CelebA dataset
(c) CelebA - Male
(d) CelebA - Smiling
@@
E.2 $\nu$ Optimization
Table 4 presents a sensitivity analysis of the degrees of freedom parameter $\nu$ for the C-$t^{3}$VAE on SVHN-LT and CIFAR100-LT, using the optimal $\beta$ from the previous study. Consistent with prior work (Kim et al., 2024), we find that $\nu$ = 10 yields robust performance on average, though slight gains can be achieved by fine-tuning within the range $[2.5,20]$. Ultimately, however, the generative FID remains relatively insensitive to variations in $\nu$, corroborating the findings of (Kim et al., 2024) regarding reconstruction quality.
@@ E.2 $\nu$ Optimization
Table 4: Variability of the FID as a function of the standard deviation $\nu$ for the C-$t^{3}$VAE model.
E.3 $\tau$ Optimization
In this section, we evaluate the effect of the $\tau$ parameter on the SVHN-LT, CIFAR100-LT and CelebA datasets for all imbalance ratios while setting $\beta$ and $\nu$ to their previously optimized values. As shown in Figure 12, the optimal $\tau$ for SVHN-LT aligns closely with our theoretical prediction. In contrast, CIFAR100-LT consistently benefits from a larger $\tau=0.4$, yielding improved FID across all imbalance settings and outperforming C-VAE. On CelebA, $\tau$ has minimal impact and the most likely value is $\tau\approx 0.3$. The analytically derived $\tau$ provides a principled initialization grounded in divergence geometry, while dataset-specific deviations reflect encoder–decoder capacity limits rather than theoretical inconsistency, highlighting the interaction between prior heaviness and representation complexity.
@@ E.3 $\tau$ Optimization
(g) CelebA Young
(h) CelebA Male
(i) CelebA Smiling
@@
Appendix F Per-Class Evaluation